I need help understanding Monomials/Polynomials/ and Factoring.
For example,
"If the length of a rectangle is 3x-2 and the width is 3x-1, what expression will represent the area of a rectangle."
If a rectangle has a width of 3x-y & length of 5x^2+y^3, what is the area?
[Factoring the Polynomials]
2x^2-8x^4
6x^2+7x-10
Please help ! Thank you

Answer by stanbon(75887)

If the length of a rectangle is 3x-2 and the width is 3x-1, what expression will represent the area of a rectangle.
----
Area = length * width = (3x-2)(3x-1) = 9x^2-9x+2
------------------------------------------------------------
If a rectangle has a width of 3x-y & length of 5x^2+y^3, what is the area?
Area = (3x-y)(5x^2+y^3) = 15x^3+3xy^3-5x^2y-y^4
---------------------------------------------------
[Factoring the Polynomials]
2x^2-8x^4 = 2x^2(1-4x^2) = 2x^2(1-2x)(1+2x)
-------
6x^2+7x-10 = 6x^2+12x-5x-10
= 6x(x+2)-5(x+2)
---
= (x+2)(6x-5)
====================
Cheers,
Stan H.
====================